Question 6 A demand for the book is Q = 250 - 20p. The cost function is c(Q) = Q?15. What is the monopolists price?
Answer #1

2-250 - २०१ २० (१ = 250 - २ P = 250 -२ २० २० 1 = २.5 - २ P= P. TP =(13-S - 3 = (२.६२ - २ up = TR वर2= 12.5x2 2=25 units. substituting into price equation P = 12.5-25 20 p=1205-0025 P = $11.25 Mops Monopolist price = $11.25